OpenStudy (anonymous):

Find the area of a rectangle that has length √3:(9) and width √3:(21)

OpenStudy (anonymous):

@AccessDenied

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

@Agent_Sniffles

OpenStudy (anonymous):

Multiply both the values using the area of a rectangle formula which is Length * Breadth. \[\sqrt{3}(9) * \sqrt{3}(21)\]

OpenStudy (anonymous):

Or if you want it in a more simplified manner, \[\sqrt{3}*\sqrt{3} * 9 * 21\] which is nothing but 3 * 9 * 21. Can you solve it now?

OpenStudy (anonymous):

thanks